MARTINA LEAVES HOME AT 9 A.M. BICYCLING AT A RATE OF 24 MI/H.
TWO HOURS LATER JOHN LEAVES DRIVING AT THE RATE OF 48 MI/H.
AT WHAT TIME WILL JOHN CATCH UP WITH MARTINA?
THANKS
JMS

Answer by checkley71(8403)

ASSUMING THEY ARE GOING IN THE SAME DIRECTION THEN:
24T=48(T-2)
24T=48T-96
24T-48T=-96
-24T=-96
T=-96/-24
T=4 HOURS AFTER 9 THEY WILL MEET.
PROOF.
24*4=48*2
96=96